OpenStudy (anonymous):

Prove that (n^2)/ (1+n^2) is decreasing, in other words, that... ((n+1)^2/(1+(n+1)^2)) is less than (n^2/(1+n^2))

OpenStudy (anonymous):

That is an increasing function f(0) = 0 f(1) = 1/2 f(2) = 4/5 etc.

OpenStudy (anonymous):

That's not a valid proof.